OpenStudy (anonymous):

A square banner has an area of 289 square inches. What is the length of one side of the banner?

OpenStudy (mathstudent55):

What is the formula for the area of a square with side L?

OpenStudy (anonymous):

it doesnt say... are you talking about the answer choices

OpenStudy (mathstudent55):

No. Let's say I give you the square in the drawing below and ask you to find its area. What formula would you use? As you know, all sides of a square have the same length, and in this case the length is L. What is the area of the square? 

OpenStudy (anonymous):

17?

OpenStudy (mathstudent55):

That's the answer to your problem.

OpenStudy (anonymous):

thanks bro your a good help

OpenStudy (mathstudent55):

wlcm \(A = L^2\) \( L = \sqrt{A} \) \( L = \sqrt{289} = 17 \)